OpenStudy (anonymous):

Use the following scenario to answer the question: A bag contains 5 green marbles, 4 red marbles, 7 blue marbles, and 9 yellow marbles. If 3 marbles are drawn from the bag one at a time but replaced each time, what is the probability of drawing a green, red and yellow marble in that order? a. 7/506 b.36/3125 c.2/3 d.1/96

OpenStudy (sloppycanada):

K so - Welcome to multiplication. We do this because the events are not reliant on each other, the marble is replaced each time. Multiply the following together - \[\frac{ 5 }{ 25 } x \frac{ 4 }{ 25 } x \frac{ 9 }{ 25 }\]

OpenStudy (anonymous):

180/15625 ?

OpenStudy (anonymous):

idk how to simplify though

OpenStudy (sloppycanada):

Find a common multiple. Can 15625 be divided by ten?

OpenStudy (sloppycanada):

Can it be divided by 5? 3?

OpenStudy (anonymous):

by 10 its 1562.5 by 5 its 3125 by 3 its 5208.3 so my guess is it would be by 5 ?

OpenStudy (sloppycanada):

Yup! Can you divide 3125 by any of the following - 6 - ? 3 - ? 2 - ? If yes, which one? If not, then 36/3125 is your answer.

OpenStudy (anonymous):

no not by any i think your right its 36/3125 !

OpenStudy (anonymous):

is there any way to simplify it agien ? i got the answer wrong :(

OpenStudy (sloppycanada):

There shouldn't be. They're all mutually exclusive events since the marble is replaced.

OpenStudy (anonymous):

alrighty then thanks ! (: